Leg A= x^2
Leg B= x^2
Hypotenuse= 8.5 in
What is x?
AKA 8.5^2=x^2+x^2
What is x?

Answer by ikleyn(52835)

.
Leg A= x^2

Leg B= x^2


Hypotenuse= 8.5 in


What is x?

AKA 8.5^2=x^2+x^2       <<<---=== If Leg A = x^2  and  Leg B = x^2   then  hypotenuse  8.5^2 = x^4 + x^4
                                  which is not what you wrote.

What is x?

--------------
Double check your writing and your thoughts.